OpenStudy (anonymous):

you invest 500 in an account with an annual interest rate of 1.1%, compounded continuously. how much money is in the account after 5 years?

OpenStudy (anonymous):

1.1% of 500 = 5.5 after 1 year = 500+5.5=505.5 1.1% of 505.5 = 5.561 after 2 years = 505.5+ 5.561 1.1% of (505.5+5.561) etc.... so this until you get five years :)

OpenStudy (anonymous):

I got 522.242 that's not any of the answers tho

OpenStudy (anonymous):

the closest one to is 528 and in the question it says round to nearest whole number would that be it?

OpenStudy (anonymous):

let me do it myself and i'll compare with you in a minute :)

OpenStudy (anonymous):

ok thanks:)

OpenStudy (anonymous):

Okay I got 528.111692 which would round down to 528 :)

OpenStudy (anonymous):

thank u so much!!

OpenStudy (anonymous):

No problem dear :)